OpenStudy (anonymous):

A system of equations is given below: y = 4x − 3 2x + 7y = 41 Which of the following steps could be used to solve by substitution? 2(y = 4x − 3) + 7y = 41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41 2(y = 4x − 3)

OpenStudy (k_lynn):

If the first equation says y=4x-3 then you would be able to put in 4x-3 for every y you see in this equation (2x+7y=41), right?

OpenStudy (anonymous):

so 4x-3=2x+7y

OpenStudy (k_lynn):

No, not quite. \[4x-3\neq2x+7y\]But\[y =4x-3\]

OpenStudy (k_lynn):

All you have to do is put in 4x-3 in for every y you see in this equation (2x+7y=41) Because \[y=4x-3\]

OpenStudy (anonymous):

so 2x + 7(4x − 3) = 41

OpenStudy (k_lynn):

Yep, you got it. \[2x+7(4x-3)=41\]

OpenStudy (anonymous):

Thank you i was really confused

OpenStudy (k_lynn):

Sure, no problem. I made a tutorial yesterday about Solving systems of equations using substitution. If you want to, you can take a look at it. http://openstudy.com/study#/updates/551aa38ae4b05a89ab81ab30

OpenStudy (anonymous):

Thank you

OpenStudy (k_lynn):

You're welcome.